The Evanec Company's next expected dividend, D1, is $3.44; its growth rate is 5%; and its common stock now sells for $30.00. New stock (external equity) can be sold to net $27.00 per share.

  1. What is Evanec's cost of retained earnings, rs? Do not round intermediate calculations. Round your answer to two decimal places.

    rs =   %

  2. What is Evanec's percentage flotation cost, F? Round your answer to two decimal places.

    F =   %

  3. What is Evanec's cost of new common stock, re? Do not round intermediate calculations. Round your answer to two decimal places.

    re =   %
Answer #1

a)

cost of retained earnings = D1/price + growth rate

= 3.44/30 + 0.05

= 16.47%

b)

flotation cost = (30-27)/30 = 3/30 = 10%

c)

cost of new common stock = D1/price + growth rate

= 3.44/27 + 0.05

= 17.74%